OpenStudy (anonymous):

A system of equations is shown below: 3x − y = 2 x + y = 6 The x-coordinate of the solution to this system of equations is _____. Numerical Answers Expected! Answer for Blank 1: ----------------

OpenStudy (error1603):

y=-5

OpenStudy (mathwizzard3):

x=11

OpenStudy (error1603):

x=1

OpenStudy (error1603):

3x+-5=-2 3x=3 x=1

OpenStudy (mathwizzard3):

1 - 5 doesn't equal 6 though

OpenStudy (mathwizzard3):

That is if you do the other equation.

OpenStudy (error1603):

I would solve for x with the first equation, then use that to solve for y in the second, then check by inputting values. 3x+y=-2 x-y=6 x=6+y

OpenStudy (anonymous):

thanks

OpenStudy (error1603):

then... 3(6+y)+y=-2 18+3y+y=-2 18+4y=-2 4y=-20 y=-5

OpenStudy (error1603):

You see?

OpenStudy (anonymous):

Yes. So y=-5 is the answer?

OpenStudy (error1603):

No now 3x+-5=-2 3x=3 x=1. x=1is the answer

OpenStudy (anonymous):

Alrighty! Thanks.

OpenStudy (error1603):

No problem